A 100g sample of Aluminum ( C= 0.90J/gC)at 120°C and a 150g sample of Iron metal (C=0.45J/C)at 90°C were placed in a container of 80g of water at 15°C. What is the final temperature of the mixture?


Can you show me the solution?


1
Expert's answer
2022-04-19T15:27:29-0400

The equation of the thermal equilibrium in this case is:


"c_am_a(T_a-T_0) + c_im_i(T_i-T_0) = c_wm_w(T_0-T_w)"

where "c_a = 0.9J\/g\/C, c_i = 0.45J\/g\/C, c_w=4.2J\/g\/C" are the specific heat capacities of the corresponding materials, "m_a = 100g, m_i = 150g, m_w=80g" are the masses, and "T_a = 120\degree C, T_i = 90\degree C, T_w=15\degree C" are the initial temperatures. "T_0" is the final temperature. Expressing "T_0", obtain:


"T_0 = \dfrac{c_am_aT_a + c_im_iT_i+c_wm_wT_w}{c_am_a + c_im_i+ c_wm_w}"

Substituting numbers, get:


"T_0 = \dfrac{0.9\cdot 100\cdot 120 +0.45\cdot 150\cdot 90+80\cdot 4.2\cdot 15}{0.9\cdot 100 +0.45\cdot 150+80\cdot 4.2} \approx 44\degree C"

Answer. "44\degree C".